Jason has 50 coins, all in 5 and 1 peso coins, amounting to Php 130.00. How many 5 peso coin does he have?

Found 3 solutions by Edwin McCravy, 12345qwerty, greenestamps:
Answer by Edwin McCravy(20054)

Jason has 50 coins, all in 5 and 1 peso coins, amounting to Php 130.00. How many 5 peso coin does he have?

Let x = number of 5 peso coins
Then 50-x = number of 1 peso coins

5x + 1(50-x) = 130

 5x + 50 - x = 130
     
     4x + 50 = 130

          4x = 80

           x = 20

He has 20 5-peso coins

and

 50-x = 

50-20 = 

30 1-peso coins

Edwin

Answer by 12345qwerty(1)

Can you answer this
A sum of money amounting to P37 consists of 1-peso bills, 2-peso bills and 5-peso bills. The number of 1-peso bills exceeds seven times the number of 5 peso bills by one, and the number of 2-peso bills is nine less than the number of 1-peso bills. How many of each kind of bills are there?

Answer by greenestamps(13198)



Algebraically....

x+y=50 the number of coins is 50
5x+1y=130 the value of the coins is 130 pesos

The difference between the two equations is 4x=80, leading to x=20; that makes y=50-20=30

ANSWER: 20 5-peso coins and 30 1-peso coins

CHECK: 20(5)+30(1)=100+30=130

Informally (if an algebraic solution is not required, and if the speed of getting the answer is important)....

(1) imagine starting with all 50 coins being 1-peso coins; that would make the total 50 pesos, which is 80 pesos short of the actual total.
(2) exchanging a 1-peso coin for a 5-peso coin keeps the total number of coins the same but increases the total value by 5-1=4 pesos.
(3) 80/4 means there have to be 20 5-peso coins; and that means 30 1-peso coins

The words of explanation make this informal solution longer than the formal algebraic solution. But without all the words, here are the complete calculations for this method:

50(1)=50
130-50=80
80/4=20
50-20=30